OpenStudy (anonymous):

Find the limit of the function by using direct substitution lim(2e^x sinx) where x-> pi/2

OpenStudy (anonymous):

Okay. just plug in pi/2 into the equation

OpenStudy (anonymous):

2e^(pi/2) * Sin(pi/2)

OpenStudy (anonymous):

(2e^x sin x) x = 2e^(pi/2)*Sin(pi/2)*(pi/2) = 2e^(pi/2)*(pi/2)

OpenStudy (anonymous):

= pi * e^(pi/2) that's the final answer

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @RobertKo

OpenStudy (anonymous):

@Tom_Boy_Rebel options are a. 0 b. 2e pi/2 c. 1 d. pi/2